### v1

@@ 4.2 Robustness Checks
We conduct a number of robustness checks and present the results in the appendix. First, we consider a number of alternative outcomes and show that our results are robust under alternative definitions of the output measure and hold for other types of GitHub actions, as well (Table B.4). [^19] Table B.4 also presents estimates on Releases, confirming the results presented in a previous version of the paper. Second, the overall changes in output quantity and quality could be driven by unobserved factors occurring at the same time as the introduction of the ban. [^20] For example, Easter Monday, an important public holiday in Italy, fell on April 10. In Section B.2, we present the results of a placebo test for a set of GitHub actions—creating a wiki page for a repository and making a repository public—for which ChatGPT is not used. Third, in Table B.5, we replace the binary outcome variables with continuous variables and show that our results are also robust on the intensive margin. Fourth, while we believe that the unexpected implementation of the ban in only one European country makes the assumption of parallel trends credible in our empirical setting, we nevertheless conduct a sensitivity analysis of possible post-treatment violations of parallel trends relative to pre-treatment violations. Appendix Figure B.2 presents relative magnitudes bounds, $\Delta^{RM}\left(\bar{M}\right)$, for the average treatment effect over the first two post-treatment days and main outcomes under different values of $\bar{M}$, as suggested by Rambachan and Roth (2023). [^21] For instance, $\bar{M}=1$ restricts the post-treatment violations of parallel trends over the first two post-treatment days to be no larger than the maximal pre-treatment violation of parallel trends over two consecutive days. Finally, for a subset of users working on multiple repositories, we construct a new panel dataset at the user–repository–day level (Appendix Section C). This allows us to include repository (project) and user fixed effects. Despite a reduction in the number of users in both the control and treatment groups, the results largely confirm the patterns from our baseline, user-level, analysis. Interestingly, we also find a small negative effect of the ban on experienced users’ output quantity relative to that of their European peers working on the same repository.
@@ 4.3 Discussion
It is possible that some Italian users immediately used tools (e.g., VPNs) to circumvent the ban. Using data on Google searches for VPN services and usage data for TOR [^22] The TOR (The Onion Router) network is an open-source overlay network of thousands of network relays that conceals a user’s IP address. Unfortunately, we cannot access actual VPN usage data at daily level. (see Appendix D), we show a sudden jump in circumvention activity among Italian Internet users in the days after the ban. Despite the easy access to circumvention technology, many corporations and organisations actually prohibit the use of VPN and TOR tools on their devices and networks, implying that their use may be limited to mainly private devices and home networks. More importantly, we still find systematic effects on output despite this circumvention activity, and one can interpret our results as a lower bound. Another concern is that our finding of heterogeneity between less experienced and experienced users could be driven by the latter’s greater skill in circumventing the ban. However, the systematic effects of the ban on tasks related to fixing bugs and closing issues suggests that this is not the case.
There are a number of follow-up questions that we cannot empirical analyse because of limitations in the data. First, while generative AI might disrupt the production flow of less experienced workers by providing faulty results, another possibility is that ChatGPT is simply a distraction. While it is not clear how this would explain the effect heterogeneity, more detailed data on the actual use of ChatGPT could help inform the design of workplace policies around generative AI. [^23] For example, existing generative AI tools such as GitHub Copilot are in general productivity enhancing because they are designed for specific tasks; Copilot, for example, only completes code and is not an open-ended chatbot. Second, more detailed data would also shed light on the question of why, after the initial increase in output and quality, we observe a decrease in the effect size for less experienced users in subsequent days. One explanation, in line with the conclusions of Kabir et al. (2024) and Li et al. (2023), could be that less experienced users still prefer to use ChatGPT as a support tool because it generates accessible and easy-to-use responses and because the costs of pursuing alternative solutions (e.g., acquiring the necessary coding skills) is relative high. Finally, our study provides evidence on the productivity effects of (the ban on) generative AI in only the very short run because the ban was short-lived and circumventing it was relatively easy.
@@ 5 Conclusion
We present novel evidence of the short-term effects of generative AI (ChatGPT) on the productivity of knowledge workers using high-frequency, observational data from over 36,000 software developers in Italy and other European countries. We use the sudden ban on ChatGPT in Italy as a natural experiment and show that the access restriction distorted output quantity and quality. Our results not only present some first empirical evidence of the widespread adoption of ChatGPT in software and code development but also show that the productivity effects of ChatGPT (and restrictions on it) differ by skill level. Our findings have the following policy implications: For some, more complex tasks, generative AI can produce faulty and erroneous output that is difficult to detect, in particular for less experienced individuals. This calls for a more targeted use of the tool in both education and work. AI-based tools that harness the power of LLMs in a more controlled form, that generate a clearly defined output and that are not based on simple text prompts (e.g., GitHub Copilot) offer guard rails to ensure more domain-specific use. Our findings also indicate that even well-intended government-mandated blocking of digital technology (to protect privacy) can lead to short-term output disruptions and costs for society. Sudden bans can be easily circumvented with VPN tools, but these adjustment activities simultaneously distort production processes and negatively impact productivity in professions that rely on the banned technology. Thus, our research also implies that policymakers should consider the potential economic cost of digital technology bans before imposing them.
@@
Appendix A Data Appendix
{ThreePartTable}{TableNotes}
[…]
Table A.1: Variable Definitions (continued)
Variable
Description
@@ Appendix A Data Appendix
Output
# Commits [PushEvent\$.size] + # Issues closed [IssuesEvent\$.action == ’closed’]+ # Pull requests closed [PullRequestEvent\$.action == ’closed’] + # Releases [ReleaseEvent]
@@
# Commits [PushEvent\$.size] + # Comments on issues [IssueCommentEvent]+ # Comments on pull requests [PullRequestReviewCommentEvent] + # Comments on commits [CommitCommentEvent] + # Pull requests [PullRequestEvent] + # Issues [IssuesEvent]
Issue closed # Issues closed [IssuesEvent\$.action == ’closed’]
Commit
# Commits [PushEvent\$.size]
Releases # Releases [ReleaseEvent]
Pull Request (PR)
# Pull requests [PullRequestEvent]
@@
Bug fix ratio (# Issues closed with “error” or “bug” label [IssuesEvent\$.action == ’closed’ & REGEXP_CONTAINS(IssuesEvent\$.label, ’error|bug’}]) / (# Issues closed [IssuesEvent\$.action == ’closed’])
PR merged
# Closed pull requests that were merged [PullRequest\$.pull_request.merged == ’true’ & PullRequestEvent\$.action == ’closed’]
PR merge ratio ( # Closed pull requests that were merged [PullRequest\$.pull_request.merged == ’true’ & PullRequestEvent\$.action == ’closed’]) / (# Pull Requests closed [PullRequestEvent\$.action == ’closed’])
C – Task Choice
Files edited per PR Average # files edited per closed pull request [AVG(PullRequestEvent\$.pull_request.changed_files) IF PullRequestEvent\$.action == ’closed’]
Lines added per PR
Average # lines added per closed pull request [AVG(PullRequestEvent\$.pull_request.additions) IF PullRequestEvent\$.action == ’closed’]
@@
Easy issue ratio
(# Easy issues opened + # Easy issues closed)/(# Issues opened + # Issues closed)
D – Placebo
Wiki pages # Wiki pages created [GollumEvent]
Public repos
# Repositories made public [PublicEvent]
@@
Appendix B GitHub User-Level Data
B.1 Descriptive Statistics
Table B.1: Descriptive Statistics
@@ B.1 Descriptive Statistics
A – User–Day Level (N = 290,992)
   Output 0.2286 0.4199 0.2319 0.4220 0.2251 0.4177 0.2644 0.4410
   Output (Shen, 2023) 0.2283 0.4198 0.2316 0.4219 0.2249 0.4175 0.2645 0.4411
   Commit 0.2215 0.4153 0.2273 0.4191 0.2155 0.4111 0.2522 0.4343
   Issue closed 0.0142 0.1183 0.0089 0.0938 0.0198 0.1394 0.0267 0.1612
   Bug fix ratio 0.0026 0.0484 0.0013 0.0350 0.0039 0.0595 0.0052 0.0689
   PR merged 0.0371 0.1890 0.0265 0.1607 0.0483 0.2144 0.0635 0.2439
   PR merge ratio 0.0396 0.2144 0.0281 0.1779 0.0518 0.2467 0.0684 0.2852
   Files edited per PR 0.0698 0.4085 0.0549 0.3777 0.0857 0.4381 0.1087 0.4780
   Lines added per PR 0.1593 0.9069 0.1259 0.8346 0.1947 0.9765 0.2480 1.0805
B – User Level (N = 36,374)
@@
   Followers 29.49 203.18 0 6 17421
   Following 19.67 185.37 0 5 28300
   Repositories 29.91 54.54 0 17 3900
@@
  • Panel A presents descriptive statistics for the baseline sample period March 27–30 (Pre) and April 3–6 (Post). The “Less experienced” sample includes all GitHub user accounts created after or in 2017 (median), while the “Experienced” sample comprises all GitHub user accounts created before 2017. The “Pkg. contributor” sample comprises all GitHub user accounts that are the owner and/or contributor to a(n) (analytical) programming package repository. The number of unique GitHub user accounts for the entire baseline sample (“Overall”) and each of the subsamples is presented in parentheses below. Panel B provides information on the individual characteristics of all GitHub user accounts in the baseline sample.
Figure B.1: Distribution of Output Quantities
@@
B.2 Additional Results
Table B.2: Effect of ChatGPT Ban on GitHub Output – March 27–28 (Pre) – April 3–4 (Post)
Output Quantity Output Quality Task Complexity
Output Output (Shen, 2023) Commit Issue closed Bug fix ratio PR merged PR merge ratio Files edited per PR Lines added per PR
(1) (2) (3) (4) (5) (6) (7) (8) (9)
Treated $\times$ Post 0.0118 0.0134 0.0111 -0.0004 0.0002 0.0064 0.0069 0.0108 0.0238
(0.0055) (0.0055) (0.0055) (0.0015) (0.0006) (0.0021) (0.0023) (0.0047) (0.0104)
Post -0.0101 -0.0105 -0.0093 -0.0007 -0.0001 -0.0025 -0.0030 -0.0052 -0.0098
(0.0024) (0.0024) (0.0024) (0.0007) (0.0003) (0.0011) (0.0012) (0.0023) (0.0052)
Overall (N = 145,496) Dep. var. mean 0.2325 0.2320 0.2251 0.0144 0.0026 0.0382 0.0406 0.0708 0.1621
Treated $\times$ Post 0.0237 0.0237 0.0221 0.0007 -0.0001 0.0099 0.0106 0.0201 0.0446
(0.0078) (0.0078) (0.0078) (0.0016) (0.0006) (0.0024) (0.0025) (0.0058) (0.0126)
Post -0.0185 -0.0184 -0.0181 -0.0009 -0.0003 -0.0035 -0.0037 -0.0092 -0.0195
(0.0035) (0.0035) (0.0035) (0.0008) (0.0003) (0.0013) (0.0014) (0.0031) (0.0068)
Less experienced (N = 74,864) Dep. var. mean 0.2355 0.2351 0.2305 0.0090 0.0015 0.0271 0.0286 0.0550 0.1264
Treated $\times$ Post -0.0010 0.0025 -0.0006 -0.0017 0.0005 0.0025 0.0026 0.0004 0.0009
(0.0078) (0.0077) (0.0076) (0.0026) (0.0012) (0.0036) (0.0041) (0.0076) (0.0170)
Post -0.0015 -0.0022 -0.0002 -0.0005 $0$0.0015 -0.0023 -0.0011 0.0002
(0.0033) (0.0033) (0.0032) (0.0011) (0.0005) (0.0017) (0.0019) (0.0035) (0.0079)
Experienced (N = 70,632) Dep. var. mean 0.2292 0.2288 0.2194 0.0201 0.0039 0.0499 0.0534 0.0876 0.1999
Treated $\times$ Post 0.0147 0.0158 0.0174 -0.0076 -0.0068 0.0006 0.0006 0.0004 -0.0130
(0.0145) (0.0144) (0.0143) (0.0054) (0.0025) (0.0074) (0.0083) (0.0153) (0.0327)
Post -0.0101 -0.0102 -0.0097 -0.0026 0.0006 0.0004 0.0007 0.0007 0.0060
(0.0060) (0.0060) (0.0059) (0.0023) (0.0010) (0.0032) (0.0037) (0.0063) (0.0143)
Pkg. contributor (N = 23,680) Dep. var. mean 0.2711 0.2710 0.2588 0.0269 0.0053 0.0677 0.0721 0.1131 0.2565

### v2

@@ 4.2 Robustness Checks
We conduct a number of robustness checks and present the results in the appendix.
Alternative outcomes:
First, we consider a number of additional outcomes in Panel A of Table LABEL:tab:did-alt-out. We show that our results are largely robust to alternative measures for output quantity (Any Event and Commit), output quality (PR merged and PR merge ratio (Holub and Thies, 2023)) as well as task complexity (Avg. files edited per merged PR). In Panel B, we additionally show that our baseline findings are qualitatively stable at the intensive margin when we use continuous outcome variables instead of binary indicators.
Placebo tests:
Second, the overall changes in output quantity and quality could be driven by unobserved factors occurring at the same time as the introduction of the ban or general differences in Italian Github user activity relative to their European peers in the work week leading up to the Easter break. To address these concerns, we undertake two placebo exercises. In the first placebo test, we assume that the ban was implemented on the weekend prior to our actual pre-treatment period from Monday 27 to Thursday 30 March 2023. The estimated placebo treatment effects are presented in Table LABEL:tab:did-placebo-2023 and are relatively small in magnitude and vastly insignificant. Importantly, we don’t find the heterogeneous effects across users uncovered in our baseline analysis. Our second placebo treatment period spans the work week prior to Good Friday on 15 April 2022. Reassuringly, our placebo estimates (Table LABEL:tab:did-placebo-2022) are again exceedingly insignificant and do not follow a clear and comparable pattern to our baseline results.
Repository-level Analysis:
For a subset of users working on multiple repositories, we construct a new panel dataset at the user–repository–day level (Section LABEL:sec:repo-results in the appendix). [^26] Note that the set of public repositories is also restricted to those with at least one user from both the control and treatment group. This allows us to include repository (project) and user fixed effects. Exploiting only within user-repository variation projects we estimate the effect of the ban on users while holding the complexity of a project constant. The results in Table LABEL:tab:did-repo-main largely confirm the patterns from our baseline, user-level, analysis. For less experienced users, we still observe a positive and significant effect on output quantity and quality, while the likelihood of closing an issue for is still negatively affected by the ban. Interestingly, we now also find some suggestive evidence for the negative impact of the ban on experienced users’ output quantity relative to that of their European peers working on the same repository. Moreover, we find additional support for the negative effect of the ban on experienced users when conditioning on programming language and restrict our attention to highly complex official package repositories (Table LABEL:tab:did-repo-lang). [^27] Because of the limited number of observations for package repositories of other programming languages, our analysis is, unfortunately, constrained to official python packages from PyPi. In addition, we exclusively present results for output quantity and quality in LABEL:tab:did-repo-lang due to the limited variation in the data for most outcomes concerning task choice and complexity.
Additional Robustness Checks
We conduct a number of additional sensitivity checks. First, we conduct a ”Leave-One-Out” analysis to further alleviate concerns that our baseline findings are dependent on our selection of control group countries. In particular, we show in Figure LABEL:fig:leave-one-out that our results are robust to excluding users from each of the control group countries. Finally, we address concerns that our results could suffer from an overrejection of the null hypotheses as a result of reuse of the identifying exogenous variation for multiple outcomes (and subsamples). Table LABEL:tab:did-rwolf reports p-values (in brackets) that are corrected for multiple hypothesis testing following the procedure described in Romano and Wolf (2005a, b, 2016). [^28] In recent work, Heath et al. (2022) show that the employed Romano–Wolf correction procedure performs well in a multitude of settings and across different dimensions. Importantly, output quantity as defined in Shen (2023) and, in particular, output quality retain their significance even after correcting for multiple hypothesis testing.
@@ 4.3 Discussion
It is possible that some Italian users immediately used tools (e.g., VPNs) to circumvent the ban. Using data on Google searches for VPN services and usage data for TOR [^29] The TOR (The Onion Router) network is an open-source overlay network of thousands of network relays that conceals a user’s IP address. Unfortunately, we cannot access actual VPN usage data at daily level. (see appendix LABEL:sec:google-tor), we show a sudden jump in circumvention activity among Italian Internet users in the days after the ban. Despite the easy access to circumvention technology, many corporations and organisations actually prohibit the use of VPN and TOR tools on their devices and networks, implying that their use may be limited to mainly private devices and home networks. More importantly, we still find systematic effects on output despite this circumvention activity, and one can interpret our results as a lower bound. Another concern is that our finding of heterogeneity between less experienced and experienced users could be driven by the latter’s greater skill in circumventing the ban. However, the systematic effects of the ban on tasks related to closing issues and the negative output effect detected in repository-level analysis suggests that this is not the case.
There are a number of follow-up questions that we are unable to empirically analyse because of limitations in the data. First, while generative AI might disrupt the production flow of less experienced workers by providing faulty results, another possibility is that ChatGPT is simply a distraction. While it is not clear how this would explain the effect heterogeneity, more detailed data on the actual use of ChatGPT could help inform the design of workplace policies around generative AI. [^30] For example, existing generative AI tools such as GitHub Copilot are in general productivity enhancing because they are designed for specific tasks; Copilot, for example, only completes code and is not an open-ended chatbot. Second, more detailed data would also shed light on the question of why, after the initial increase in output and quality, we observe a decrease in the effect size for less experienced users in subsequent days. One explanation, in line with the conclusions of Kabir et al. (2024) and Li et al. (2023), could be that less experienced users still prefer to use ChatGPT as a support tool because it generates accessible and easy-to-use responses and because the costs of pursuing alternative solutions (e.g., acquiring the necessary coding skills) is relatively high. Finally, our study provides evidence on the productivity effects of (the ban on) generative AI in only the very short run because the ban was short-lived and circumventing it was relatively easy.
@@ 5 Conclusion
We present novel evidence of the short-term effects of generative AI (ChatGPT) on the productivity of knowledge workers using high-frequency, observational data from over 36,000 software developers in Italy and other European countries. We use the sudden ban on ChatGPT in Italy as a natural experiment and show that the access restriction distorted output quantity and quality. Our results not only present some first empirical evidence of the widespread adoption of ChatGPT in software and code development but also show that the productivity effects of ChatGPT (and restrictions on it) differ by experience level. Our findings have the following policy implications: For some, more complex tasks, generative AI can produce faulty and erroneous output that is difficult to detect, in particular for less experienced individuals. This calls for a more targeted use of the tool in both education and work. AI-based tools that harness the power of LLMs in a more controlled form, that generate a clearly defined output and that are not based on simple text prompts (e.g., GitHub Copilot) offer guard rails to ensure more domain-specific use. Our findings also indicate that even well-intended government-mandated blocking of digital technology (to protect privacy) can lead to short-term output disruptions and costs for society. Sudden bans can be easily circumvented with VPN tools, but these adjustment activities simultaneously distort production processes and negatively impact productivity in professions that rely on the banned technology. Thus, our research also implies that policymakers should consider the potential economic cost of digital technology bans before imposing them.
A: Output Quantity and Quality
Output Output (Shen, 2023) Output (Holub & Thies, 2023) Issue closed PR merge ratio
Overall Treated $\times$ Post 0.0062 0.0102 0.0090 -0.0045** 0.0045
(N = 290,864) (0.0076) (0.0075) (0.0080) (0.0022) (0.0030)
Dep. var. mean 0.2314 0.2283 0.2624 0.0142 0.0359
Less experienced Treated $\times$ Post 0.0193* 0.0216** 0.0203* -0.0017 0.0118***
(N = 149,680) (0.0107) (0.0107) (0.0111) (0.0023) (0.0035)
Dep. var. mean 0.2339 0.2316 0.2519 0.0089 0.0258
Experienced Treated $\times$ Post -0.0079 -0.0020 -0.0030 -0.0077* -0.0034
(N = 141,184) (0.0107) (0.0106) (0.0116) (0.0039) (0.0052)
Dep. var. mean 0.2287 0.2249 0.2734 0.0198 0.0467
Pkg. contributor Treated $\times$ Post 0.0181 0.0164 0.0270 -0.0151* 0.0113
(N = 47,328) (0.0205) (0.0200) (0.0215) (0.0084) (0.0104)
Dep. var. mean 0.2688 0.2644 0.3312 0.0267 0.0615
B: Task Choice and Complexity
PR opened Avg. lines added per opened PR Avg. lines added per merged PR Easy issue closed Interactive Activity
Overall Treated $\times$ Post $0$0.0039 0.0220 -0.0004 0.0006
(N = 290,864) (0.0034) (0.0148) (0.0135) (0.0004) (0.0043)
Dep. var. mean 0.0402 0.1581 0.1439 0.0007 0.0637
Less experienced Treated $\times$ Post 0.0040 0.0211 0.0456*** -0.0004 0.0037
(N = 149,680) (0.0040) (0.0186) (0.0164) (0.0004) (0.0045)
Dep. var. mean 0.0315 0.1358 0.1148 0.0005 0.0340
Experienced Treated $\times$ Post -0.0046 -0.0324 -0.0031 -0.0003 -0.0028
(N = 141,184) (0.0056) (0.0237) (0.0221) (0.0008) (0.0077)
Dep. var. mean 0.0494 0.1817 0.1749 0.0009 0.0953
Pkg. contributor Treated $\times$ Post -0.0221* -0.0933* 0.0226 -0.0021* 0.0127
(N = 47,328) (0.0117) (0.0477) (0.0422) (0.0011) (0.0158)
Dep. var. mean 0.0661 0.2356 0.2210 0.0014 0.1393
  • Notes:
  • All specifications include user–fixed effects, day-of-the-week–fixed effects, and a linear time trend for the control and treatment group. The “Less experienced” sample includes all Github user accounts created after or in 2017 (median), while the “Experienced” sample comprises all GitHub user accounts created before 2017. The “Pkg. contributor” sample comprises all GitHub user accounts that are the owner and/or contributor to a (analytical) programming package repository. The number of observations is depicted in parentheses after each sample definition. A log plus one transformation is applied to Avg. lines added per PR (opened or merged). Robust standard errors in parentheses are clustered on the user-level: * $p<0.1$, ** $p<0.05$, *** $p<0.01$.
Figure 1: Less Experienced Users – Event-Study Estimates
Notes: Event-study estimates across outcomes for “less experienced” GitHub user accounts (created after or in 2017). The sample period spans March 27–30 (Pre) and April 3–6 (Post). All specifications include user, time, and day-of-the-week fixed effects. A log plus one transformation is applied to Avg. lines added per PR (opened or merged). 95% (90%) confidence intervals for robust standard errors clustered at the user level are depicted in light (dark) grey.
@@
Appendix A Timeline
Figure A.1 depicts the timing of the events. The ban was introduced on Friday March 31st 2023. In our baseline analysis, we consider the 4 business days prior to the ban, (MON – THU, March 27th – 30th) as the pre-period. To make the analysis, comparable, we ignore the two days following the introduction of the ban which were weekend days (SAT and SUN, April 1st – 2nd, in gray are excluded). The Post period is defined as the first 4 business days, after the introduction of the ban (MON – THU, April 3rd – 6th 2023). We further exclude Friday, 7th 2023, because this was Good Friday, a national holiday in all European countries under consideration in our sample.
Figure A.1: Time Line
PRE PeriodBan Intro/WE (Exc.)POST PeriodEaster (Exc.)MON 27/03THU 30/03FRI 31/03SUN 02/04MON 03/04THU 06/04
Appendix B Data Appendix
B.1 Github User Location
Figure B.1: GitHub User Profile
Notes: Profile last accessed on 07 May 2024 at https://github.com/0x0f0f0f.
We scrape, among others, the following key attributes at the user-level for a defined set of locations (location) and span of time-period (i.e. createdAt from 1 Jan 2009 until 11 Apr 2023) from the GitHub GraphQL API: [^31] For more details on the user-level attributes please refer to the official GitHub documentation: https://docs.github.com/en/graphql/reference/objects#user.
  • login: The user’s username (e.g. 0x0f0f0f).
  • location: The user’s location (e.g. Italy).
  • followers: The number of followers the user has (e.g. 165).
  • following: The numbers of users the user is following (e.g. 88).
  • repositories: A list of user’s (public) repositories (e.g. 44).
  • createdAt: The date and time the user’s account was created.
For more details on the scraping procedure (e.g. the set of locations) and the python implementation please refer to the public GitHub repository https://GitHub.com/sodalabsio/GitHub_scrape.
B.2 Definition and Construction of Outcome Variables
{ThreePartTable}
Table B.1: Variable Definitions
Variable Description
@@ B.2 Definition and Construction of Outcome Variables
Output # Commits [PushEvent\$.size] + # Issues closed [IssuesEvent\$.action == ’closed’]+ # Pull requests closed [PullRequestEvent\$.action == ’closed’] + # Releases [ReleaseEvent]
@@
Issues closed # Issues closed [IssuesEvent\$.action == ’closed’]
Commits # Commits [PushEvent\$.size]
@@
Pull Requests (PRs) merged # Closed pull requests that were merged [PullRequestEvent\$.pull_request.merged == ’true’ & PullRequestEvent\$.action == ’closed’]
PR merge ratio ( # Closed pull requests that were merged [PullRequestEvent\$.pull_request.merged == ’true’ & PullRequestEvent\$.action == ’closed’]) / (# Pull Requests closed [PullRequestEvent\$.action == ’closed’])
PR merge ratio (Holub and Thies, 2023) ( # Opened pull requests that were merged [PullRequest\$.pull_request.merged == ’true’ & PullRequestEvent\$.action == ’closed’ & PullRequestEvent\$.action == ’opened’]) / (# Pull Requests opened [PullRequestEvent\$.action == ’opened’])
C – Task Choice and Complexity
Pull Requests (PRs) opened # Pull Requests opened [PullRequestEvent\$.action == ’opened’]
Avg. files edited per merged PR Average # files edited per merged pull request [AVG(PullRequestEvent\$.pull_request.changed_files) IF PullRequestEvent\$.pull_request.merged == ’true’ & PullRequestEvent\$.action == ’closed’]
Avg. Lines added per merged PR Average # lines added per merged pull request [AVG(PullRequestEvent\$.pull_request.additions) IF PullRequestEvent\$.pull_request.merged == ’true’ & PullRequestEvent\$.action == ’closed’]
Avg. Lines added per opened PR Average # lines added per opened pull request [AVG(PullRequestEvent\$.pull_request.additions) IF PullRequestEvent\$.action == ’opened’]
Easy issue closed # Easy issues closed [IssuesEvent\$.action == ’closed’]
@@
D – User Activity
“Work” hours Time difference (in hours) between first and last activity (any or “prodctive”)]
First activity Clock Hour of first activity (any or “prodctive”)
Last activity Clock Hour of last activity (any or “prodctive”)
Unusual activity (10th %tile) Dummy = 1 if activity (any or “prodctive”) before 6am (10th percentile of First activity) or after 9pm (90th percentile of Last activity)
Unusual activity (25th %tile) Dummy = 1 if activity (any or “prodctive”) before 8am (25th percentile of First activity) or after 7pm (75th percentile of Last activity)
@@
Appendix C GitHub User-Level Data
C.1 Descriptive Statistics
@@
A - User-Day–Level (N = 290,864)
     Output 0.2314 0.4217 0.2339 0.4233 0.2287 0.4200 0.2688 0.4433
     Output (Shen, 2023) 0.2283 0.4198 0.2316 0.4218 0.2249 0.4175 0.2644 0.4410
     Output (Holub and Thies, 2023) 0.2624 0.4399 0.2519 0.4341 0.2734 0.4457 0.3312 0.4706
     Issue closed 0.0142 0.1183 0.0089 0.0938 0.0198 0.1394 0.0267 0.1611
     PR merge ratio 0.0359 0.1846 0.0258 0.1572 0.0467 0.2092 0.0615 0.2381
     PR opened 0.0402 0.1964 0.0315 0.1746 0.0494 0.2168 0.0661 0.2484
     Avg. lines added per opened PR 0.1581 0.9003 0.1358 0.8648 0.1817 0.9359 0.2356 1.0474
     Avg. lines added per merged PR 0.1439 0.8563 0.1148 0.7928 0.1749 0.9179 0.2210 1.0124
     Easy issue closed 0.0007 0.0263 0.0005 0.0219 0.0009 0.0302 0.0014 0.0367
     Interactive Activity 0.0637 0.2442 0.0340 0.1811 0.0953 0.2936 0.1393 0.3463
B - User–Level (N = 36,358)
@@
     Followers 29.49 203.22 0 6 17421
     Following 19.67 185.41 0 5 28300
     Repositories 29.90 54.53 0 17 3900
@@
  • Panel A presents descriptive statistics for the baseline sample period Pre 27-30.03 – Post 03-06.04. The “Less experienced” sample includes all GitHub user accounts created after or in 2017 (median), while the “Experienced” sample comprises all GitHub user accounts created before 2017. The “Pkg. contributor” sample comprises all GitHub user accounts that are the owner and/or contributor to a (analytical) programming package repository. The number of unique GitHub user accounts for the entire baseline sample (“Overall”) and each of the subsamples is presented in parentheses below. A log plus one transformation is applied to Avg. lines added per PR (opened or merged). Panel B provides information on the individual characteristics of all GitHub user accounts in the baseline sample.
@@
Notes: The distribution of (any) activity by GitHub users across hours of the day is presented separately for the four days pre- and post-ban.
Notes: The distribution of (any) activity by GitHub users across hours of the day is presented separately for the four days pre- and post-ban.
C.2 Additional Results
(1) (2) (3) (4) (5)
A: Output Quantity and Quality
Output Output (Shen, 2023) Output (Holub & Thies, 2023 Issue closed PR merge ratio
Overall Treated $\times$ Post 0.0116*** 0.0128*** 0.0147*** -0.0005 0.0056***
(N = 290,864) (0.0043) (0.0042) (0.0045) (0.0011) (0.0015)
Dep. var. mean 0.2314 0.2283 0.2624 0.0142 0.0359
Less experienced Treated $\times$ Post 0.0177*** 0.0177*** 0.0186*** 0.0027** 0.0095***
(N = 149,680) (0.0061) (0.0061) (0.0062) (0.0013) (0.0018)
Dep. var. mean 0.2339 0.2316 0.2519 0.0089 0.0258
Experienced Treated $\times$ Post 0.0052 0.0079 0.0108* -0.0043** 0.0013
(N = 141,184) (0.0059) (0.0059) (0.0064) (0.0020) (0.0025)
Dep. var. mean 0.2287 0.2249 0.2734 0.0198 0.0467
Pkg. contributor Treated $\times$ Post 0.0148 0.0150 0.0194* -0.0095** -0.0001
(N = 47,328) (0.0109) (0.0108) (0.0115) (0.0040) (0.0049)
Dep. var. mean 0.2688 0.2644 0.3312 0.0267 0.0615
B: Task Choice and Complexity
PR opened Avg. lines added per opened PR Avg. lines added per merged PR Easy issue closed Interactive Activity
Overall Treated $\times$ Post 0.0035** 0.0180** 0.0249*** 0.0002 0.0011
(0.0017) (0.0078) (0.0070) (0.0002) (0.0022)
(N = 290,864) Dep. var. mean 0.0402 0.1581 0.1439 0.0007 0.0637
Less experienced Treated $\times$ Post 0.0056*** 0.0322*** 0.0402*** 0.0002 0.0016
(N = 149,680) (0.0021) (0.0100) (0.0087) (0.0002) (0.0024)
Dep. var. mean 0.0315 0.1358 0.1148 0.0005 0.0340
Experienced Treated $\times$ Post 0.0012 0.0025 0.0079 0.0001 0.0004
(N = 141,184) (0.0029) (0.0121) (0.0112) (0.0004) (0.0039)
Dep. var. mean 0.0494 0.1817 0.1749 0.0009 0.0953
Pkg. contributor Treated $\times$ Post -0.0083 -0.0269 -0.0006 -0.0010* 0.0005
(N = 47,328) (0.0057) (0.0240) (0.0219) (0.0006) (0.0081)
Dep. var. mean 0.0661 0.2356 0.2210 0.0014 0.1393